OpenStudy (anonymous):

(F of G)(x)=-7x+52 what is the domain??

OpenStudy (kira_yamato):

\[(-\infty , \infty)\]

OpenStudy (anonymous):

depends on the domain of G

OpenStudy (anonymous):

g(x)= sqrt(49-7x)

OpenStudy (kira_yamato):

\[fg(x) = -7\sqrt{49-7x}+52\] So the domain of fg(x) is \[D_{fg(x)} = [7,\infty)\]

OpenStudy (anonymous):

f(x)=x^2+3 g(x)=sqrt(49-7x) (F of G)(x)=-7x+52 What is Domain of (f of G)(X)?? Not (-inf, inf) or [7, inf)